OpenStudy (anonymous):

A certain drug is made from only two ingredients: compound A and compound B. There are 8 milliliters of compound A used for every 5 milliliters of compound B. If a chemist wants to make 1079 milliliters of the drug, how many milliliters of compound A are needed?

OpenStudy (anonymous):

Well, this is a ratio problem: First you know that the ratio of A:B is 8:5, and 8+5 = 13. Now 1079/13 = 83. That is the number of rations of each chemical you have. To get chemical A, multiply 83 with 8, and for B, multiply 5 with 83. You can check by making sure A + b = 1079.